I'm tring to get thru college math and is having a hard time understanding equations and inequalities please help
one of the problems is 8x-(3x+2)=3x-10

Answer by prabhjyot(164)

8x-(3x+2)=3x-10
first open the brackets
8x - 3x -2=3x-10
(8x-3x-3x)=-10+2
8x=-8
x=- 8/8
x=-1